### v1

@@
This paper develops a step-by-step guide to leniency designs, drawing on recent econometric literatures. The starting point of this operator’s manual is a link between leniency designs and more conventional instrumental variable regressions, in which indicators for decision-maker assignment are used directly as instruments. This link follows from observing that, with no controls in this more conventional specification, the population first-stage regression of treatment on the assignment indicators gives the ideal leniency measure. Hence, the more conventional specification can be understood as also using a leniency measure as the instrument. Complementing a recent review by [78], we use this link to to develop answers to the above questions, via a robust way of estimating the more conventional instrumental variable regression.
@@
Another approach is to bias-correct the 2SLS formula, which replaces the infeasible instrument $\tilde{\ell}_{i}$ in (3) with the estimated relative leniency $\hat{\ell}_{i}$. When 2SLS does this, it increases the expectation of the denominator from $\sum_{i}\tilde{\ell}_{i}^{2}$ to $\sum_{i}\tilde{\ell}_{i}^{2}+K\operatorname{var}(\eta_{i})$ under homoskedasticity. The quantity $\sum_{i}\tilde{\ell}_{i}^{2}$ corresponds to the numerator of the population partial R-squared statistic; it measures the increase in the explained sum of squares from adding instruments to the first stage. Since 2SLS uses in-sample fitted values to estimate the predictive power of the instruments, it overstates this predictive power, exactly analogous to how the unadjusted R-squared overstates the predictive power of regression (the same issue happens in the numerator, and taking the ratio gives the rule-of-thumb bias formula for 2SLS). Paralleling how the adjusted R-squared fixes the issue by using a degrees of freedom correction, we can use degrees of freedom adjustments in both the denominator and the numerator of the 2SLS formula. Unfortunately, similar to the adjusted R-squared formula, the resulting bias-corrected 2SLS estimator (dating back to [105]) works only under homoskedasticity.
@@
Unfortunately, with many instruments, the same issue that causes bias in the 2SLS estimator also causes its standard errors to be small relative to the infeasible estimator. Recall that overfitting inflates the denominator in eq.˜3 from $\sum_{i}\tilde{\ell}_{i}^{2}$ to $\sum_{i}\tilde{\ell}_{i}^{2}+K\operatorname{var}(\eta_{i})$ under homoskedasticity. Since the same denominator appears in the standard error formula, overfitting shrinks standard errors too. Thus, 2SLS mimics not only the ordinary least squares estimate, but also its precision. As [74] emphasize, this can mask the weak instrument problem: researchers see tight standard errors around an estimate close to ordinary least squares and may wrongly conclude the causal effect is precisely estimated with minimal selection bias.
@@ Appendix A Derivations
To derive the bias expressions for the other estimators, it is convenient to use matrix notation. We denote the vectors of $n$ treatment and outcome observations by $x$ and $y$, respectfully, and denote the matrix of covariates and instruments with rows ${w}_{i}^{\prime}$ and $z_{i}^{\prime}$ by $W$ and $Z$, respectively. To state the results, it will be convenient to link the sum of squares of the relative leniency, $\sum_{i}\tilde{\ell}_{i}^{2}$ to the first-stage $F$ statistic and the partial $R^{2}$. Recall that the first-stage $F$ statistic for testing the null that $\pi=0$ is, under homoskedastic errors, given by $F=\hat{\pi}^{\prime}\tilde{Z}^{\prime}\tilde{Z}\hat{\pi}/K\operatorname{var}(\eta_{i})$, with expectation $E[F]=\sum_{i}\tilde{\ell}_{i}^{2}/K\operatorname{var}(\eta_{i})+1$. The population partial $R^{2}$ from adding the instruments to the first-stage regression is given by $R^{2}=(\sum_{i}\tilde{\ell}_{i}^{2})/(\sum_{i}\tilde{\ell}_{i}^{2}+n\operatorname{var}(\eta_{i}))$ so
$\frac{\sum_{i}\tilde{\ell}_{i}^{2}}{\operatorname{var}(\eta_{i})}=K(E[F]-1)=\frac{nR^{2}}{1-R^{2}}.$ (6)
@@
Consider an estimator $\hat{\beta}_{G}=y^{\prime}Gx/x^{\prime}Gx$ based on a relative leniency measure $Gx$ such that $Gx=\tilde{\ell}+G\nu$. Then, under homoskedastic errors, $E[\hat{\beta}_{G}]-\beta\approx\frac{\operatorname{tr}(G)\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\sum_{i}\tilde{\ell}_{i}^{2}+\operatorname{tr}(G)\operatorname{var}(\eta_{i})}=\frac{\operatorname{tr}(G)}{K(E[F]-1)+\operatorname{tr}(G)}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\eta_{i})}$.
@@
$E[\hat{\beta}_{OLS}]-\beta\approx\frac{1-L/n}{K(E[F]-1)/n+1-L/n}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\eta_{i})}\approx(1-R^{2})\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\eta_{i})},$
@@
$E[\hat{\beta}_{TSLS}]-\beta\approx\frac{1}{E[F]}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\eta_{i})}$
@@
$E[\hat{\beta}_{JIVE}-\beta]\approx-\frac{L}{K(E[F]-1)-L}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\eta_{i})},$
@@
$\begin{pmatrix}y^{\prime}G_{UJIVE}x\\ x^{\prime}G_{UJIVE}x\end{pmatrix}=\sum_{i}\begin{pmatrix}\tilde{\ell}_{i}y_{i}\\ \tilde{\ell}_{i}x_{i}\end{pmatrix}+\sum_{i}\begin{pmatrix}r_{Yi}\nu_{i}\\ r_{i}\nu_{i}\end{pmatrix}+\sum_{i,j}\begin{pmatrix}G_{ij}\nu_{Yi}\nu_{j}\\ G_{ij}\nu_{i}\nu_{j}\\ \end{pmatrix}$,
where $G_{ij}$ is the $(i,j)$ element of $G_{UJIVE}$, while $r_{Yi}=\sum_{j}G_{ji}(z_{j}^{\prime}\pi_{Y}+w_{j}^{\prime}\delta_{Y})$ and $r_{i}=\sum_{j}G_{ji}(z_{j}^{\prime}\pi+w_{j}^{\prime}\delta)$ denote the “signal” in the reduced form and the first stage. The second and third terms represent additional noise components in the estimator relative to the infeasible estimator $\hat{\beta}^{*}$. Since the third term is uncorrelated with the first two, it follows that the covariance matrix of the left-hand side may be written
$\Sigma=\sum_{i}\operatorname{var}\begin{pmatrix}\tilde{\ell}_{i}y_{i}+r_{Yi}\nu_{i}\\ \tilde{\ell}_{i}x_{i}+r_{i}\nu_{i}\end{pmatrix}+\sum_{i,j}\begin{pmatrix}G_{ij}^{2}\sigma^{2}_{y,i}\sigma^{2}_{x,j}+G_{ij}G_{ji}\sigma_{xy,i}\sigma_{xy,j}&G_{ij}(G_{ij}+G_{ji})\sigma_{yx,i}\sigma^{2}_{x,j}\\ G_{ij}(G_{ij}+G_{ji})\sigma_{yx,i}\sigma^{2}_{x,j}&G_{ij}(G_{ij}+G_{ji})\sigma^{2}_{x,i}\sigma^{2}_{x,j}\end{pmatrix}$
where $\sigma^{2}_{y,i}=\operatorname{var}(\eta_{Y,i})$, $\sigma^{2}_{x,i}=\operatorname{var}(\eta_{i})$, and $\sigma_{yx,i}=\operatorname{cov}(\eta_{Y,i},\eta_{i})$. Furthermore, the numerator and denominator can be shown to be asymptotically normal by a martingale central limit theorem [86, Lemma D.5] so that

### v2

@@
This paper develops a step-by-step guide to leniency designs, drawing on recent econometric literatures. The starting point of this operator’s manual is a link between leniency designs and more conventional instrumental variable regressions, in which indicators for decision-maker assignment are used directly as instruments. This link follows from observing that, with no controls in this more conventional specification, the population first-stage regression of treatment on the assignment indicators gives the ideal leniency measure. Hence, the more conventional specification can be understood as also using a leniency measure as the instrument. Complementing a recent review by [78], we use this link to to develop answers to the above questions via a robust way of estimating the more conventional instrumental variable regression.
@@
Another approach is to bias-correct the 2SLS formula, which replaces the infeasible instrument $\tilde{\ell}_{i}$ in (3) with the estimated relative leniency $\hat{\ell}_{i}$. When 2SLS does this, it increases the expectation of the denominator from $\sum_{i}\tilde{\ell}_{i}^{2}$ to $\sum_{i}\tilde{\ell}_{i}^{2}+K\operatorname{var}(\nu_{i})$ under homoskedasticity. The quantity $\sum_{i}\tilde{\ell}_{i}^{2}$ corresponds to the numerator of the population partial R-squared statistic; it measures the increase in the explained sum of squares from adding instruments to the first stage. Since 2SLS uses in-sample fitted values to estimate the predictive power of the instruments, it overstates this predictive power, exactly analogous to how the unadjusted R-squared overstates the predictive power of regression (the same issue happens in the numerator, and taking the ratio gives the rule-of-thumb bias formula for 2SLS). Paralleling how the adjusted R-squared fixes the issue by using a degrees of freedom correction, we can use degrees of freedom adjustments in both the denominator and the numerator of the 2SLS formula. Unfortunately, similar to the adjusted R-squared formula, the resulting bias-corrected 2SLS estimator (dating back to [105]) works only under homoskedasticity.
@@
Unfortunately, with many instruments, the same issue that causes bias in the 2SLS estimator also causes its standard errors to be small relative to the infeasible estimator. Recall that overfitting inflates the denominator in eq.˜3 from $\sum_{i}\tilde{\ell}_{i}^{2}$ to $\sum_{i}\tilde{\ell}_{i}^{2}+K\operatorname{var}(\nu_{i})$ under homoskedasticity. Since the same denominator appears in the standard error formula, overfitting shrinks standard errors too. Thus, 2SLS mimics not only the ordinary least squares estimate, but also its precision. As [74] emphasize, this can mask the weak instrument problem: researchers see tight standard errors around an estimate close to ordinary least squares and may wrongly conclude the causal effect is precisely estimated with minimal selection bias.
@@ Appendix A Derivations
To derive the bias expressions for the other estimators, it is convenient to use matrix notation. We denote the vectors of $n$ treatment and outcome observations by $x$ and $y$, respectfully, and denote the matrix of covariates and instruments with rows ${w}_{i}^{\prime}$ and $z_{i}^{\prime}$ by $W$ and $Z$, respectively. To state the results, it will be convenient to link the sum of squares of the relative leniency, $\sum_{i}\tilde{\ell}_{i}^{2}$ to the first-stage $F$ statistic and the partial $R^{2}$. Recall that the first-stage $F$ statistic for testing the null that $\pi=0$ is, under homoskedastic errors, given by $F=\hat{\pi}^{\prime}\tilde{Z}^{\prime}\tilde{Z}\hat{\pi}/K\operatorname{var}(\nu_{i})$, with expectation $E[F]=\sum_{i}\tilde{\ell}_{i}^{2}/K\operatorname{var}(\nu_{i})+1$. The population partial $R^{2}$ from adding the instruments to the first-stage regression is given by $R^{2}=(\sum_{i}\tilde{\ell}_{i}^{2})/(\sum_{i}\tilde{\ell}_{i}^{2}+n\operatorname{var}(\nu_{i}))$ so
$\frac{\sum_{i}\tilde{\ell}_{i}^{2}}{\operatorname{var}(\nu_{i})}=K(E[F]-1)=\frac{nR^{2}}{1-R^{2}}.$ (6)
@@
Consider an estimator $\hat{\beta}_{G}=y^{\prime}Gx/x^{\prime}Gx$ based on a relative leniency measure $Gx$ such that $Gx=\tilde{\ell}+G\nu$. Then, under homoskedastic errors, $E[\hat{\beta}_{G}]-\beta\approx\frac{\operatorname{tr}(G)\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\sum_{i}\tilde{\ell}_{i}^{2}+\operatorname{tr}(G)\operatorname{var}(\nu_{i})}=\frac{\operatorname{tr}(G)}{K(E[F]-1)+\operatorname{tr}(G)}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\nu_{i})}$.
@@
$E[\hat{\beta}_{OLS}]-\beta\approx\frac{1-L/n}{K(E[F]-1)/n+1-L/n}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\nu_{i})}\approx(1-R^{2})\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\nu_{i})},$
@@
$E[\hat{\beta}_{TSLS}]-\beta\approx\frac{1}{E[F]}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\nu_{i})}$
@@
$E[\hat{\beta}_{JIVE}-\beta]\approx-\frac{L}{K(E[F]-1)-L}\frac{\operatorname{cov}(\varepsilon_{i},\nu_{i})}{\operatorname{var}(\nu_{i})},$
@@
$\begin{pmatrix}y^{\prime}G_{UJIVE}x\\ x^{\prime}G_{UJIVE}x\end{pmatrix}=\sum_{i}\begin{pmatrix}\tilde{\ell}_{i}y_{i}\\ \tilde{\ell}_{i}x_{i}\end{pmatrix}+\sum_{i}\begin{pmatrix}r_{Yi}\nu_{i}\\ r_{i}\nu_{i}\end{pmatrix}+\sum_{i,j}\begin{pmatrix}G_{ij}\nu_{Y,i}\nu_{j}\\ G_{ij}\nu_{i}\nu_{j}\\ \end{pmatrix}$,
where $G_{ij}$ is the $(i,j)$ element of $G_{UJIVE}$, while $r_{Y,i}=\sum_{j}G_{ji}(z_{j}^{\prime}\pi_{Y}+w_{j}^{\prime}\delta_{Y})$ and $r_{i}=\sum_{j}G_{ji}(z_{j}^{\prime}\pi+w_{j}^{\prime}\delta)$ denote the “signal” in the reduced form and the first stage. The second and third terms represent additional noise components in the estimator relative to the infeasible estimator $\hat{\beta}^{*}$. Since the third term is uncorrelated with the first two, it follows that the covariance matrix of the left-hand side may be written
$\Sigma=\sum_{i}\operatorname{var}\begin{pmatrix}\tilde{\ell}_{i}y_{i}+r_{Y,i}\nu_{Y,i}\\ \tilde{\ell}_{i}x_{i}+r_{i}\nu_{i}\end{pmatrix}+\sum_{i,j}\begin{pmatrix}G_{ij}^{2}\sigma^{2}_{y,i}\sigma^{2}_{x,j}+G_{ij}G_{ji}\sigma_{xy,i}\sigma_{xy,j}&G_{ij}(G_{ij}+G_{ji})\sigma_{yx,i}\sigma^{2}_{x,j}\\ G_{ij}(G_{ij}+G_{ji})\sigma_{yx,i}\sigma^{2}_{x,j}&G_{ij}(G_{ij}+G_{ji})\sigma^{2}_{x,i}\sigma^{2}_{x,j}\end{pmatrix}$
where $\sigma^{2}_{y,i}=\operatorname{var}(\nu_{Y,i})$, $\sigma^{2}_{x,i}=\operatorname{var}(\nu_{i})$, and $\sigma_{yx,i}=\operatorname{cov}(\nu_{Y,i},\nu_{i})$. Furthermore, the numerator and denominator can be shown to be asymptotically normal by a martingale central limit theorem [86, Lemma D.5] so that